OpenStudy (anonymous):

What is the solution to the following equation? 3(7x − 12) + 26 = 3x − 10 0 1 5 8

OpenStudy (anonymous):

I think it is 5

OpenStudy (anonymous):

Its 0. I used an free inline algebra solver. Would you like a link? It's quite useful

OpenStudy (haseeb96):

21x -36 +26=3x-10 18x -10= -10 18x = -10 +10 18x=0 x=0 this is your answer

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

I would like the link

OpenStudy (anonymous):

@ANA789

OpenStudy (anonymous):

http://www.marblesoftwaresolutions.com/AlgebraSolver/ there you go :)

OpenStudy (anonymous):

It dont work when I click solve

OpenStudy (anonymous):

That's odd. Can you screenshoot it?

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

Do you got a different algebra solver?

OpenStudy (anonymous):

http://www.cymath.com/ try this one